# Zonal vs. Nodal Pricing: An Analysis of Different Pricing Rules in the German Day-Ahead Market

### v4

@@ 4.1.1 Welfare losses with Euphemia - Impact of Non-Uniform Pricing
Euphemia maximizes welfare subject to linear and anonymous prices, which is different from unconstrained welfare maximization. In our analysis, the average daily welfare loss associated with Euphemia compared to the national configuration (without any network constraints) is 0.34% – or EUR 109,801.24 in absolute terms. This is comparable with previous estimate of 0.05% as a relative loss (Meeus et al.,, 2009). Disregarding market coupling, the cuts introduced by the Euphemia algorithm to get to linear prices do not deteriorate the outcome by a lot. After redispatch, Euphemia suffers an average daily loss of 9.2% (or kEUR 3,683.24) in total costs compared to the nodal allocation rule. We argue that the national, zonal, or Euphemia models differ only marginally in terms of allocation, and only a transition to a nodal allocation rule would substantially affect the average generation and redispatch costs.
@@ 4.2 Price Levels
Table 3 provides average prices over all hours of the weeks under consideration, while Table 4 illustrates median prices and Table 5 price standard deviations for all computed prices under the respective allocation and pricing rule.
@@
IP 48.66 48.44 48.84 48.23 45.79 48.66
CH 31.09 31.17 31.27 31.11 38.53 49.41
Join 48.68 49.37 48.60 48.31 49.27 49.19
Euphemia 48.84
Table 4: Median Prices
in EUR/MWh National 2 Zones (k) 2 Zones (s) 3 Zones 4 Zones Nodal
@@
Average CH prices are slightly lower compared to the other pricing rules under consideration, with lower standard deviation and irrespective of the allocation rule. In contrast, median prices are at comparable levels across all pricing rules. Since CH prices are unaffected by pre-determined generator commitments, they tend to be lower and more stable. In particular, prices can be set by a cheaper – yet uncommitted – generator (Schiro et al.,, 2016). In contrast, average Euphemia prices tend to be slightly higher than zonal prices, while median prices are in the same ranges. This finding implies that the price-setting generator is usually identical for Euphemia and zonal prices, yet sometimes the Euphemia algorithm introduces a cut that disallows a generator that would otherwise be dispatched. As a result, a higher-price generator may set the price.
@@
Under nodal pricing (ca. 40,000 prices per day for 1670 nodes – compared to, e.g., 24 national prices per day) more price outliers were observed than under zonal pricing. For instance, 0.13% of IP prices had to be capped 100 EUR/MWh under the zonal configurations, compared to 0.25% under nodal pricing. Without capping, the maximum observed IP and Join price would be as high as 62,000 EUR/MWh for an hour of extreme scarcity. Note that SDAC also imposes a price cap of 5,000 EUR/MWh for such periods. CH prices, in contrast, are less prone to outliers, peaking at 423 EUR/MWh. Thus tight flow constraints can increase price levels in the short run. The zonal allocation rules oversimplify the transmission network to an extent that nodal prices can hardly be lower, even without congestion. In contrast, when congestion is present in parts of the grid, nodal prices will be higher, resulting in higher average and median prices. Upon further analysis, it was found that nodal prices have high sensitivity to transmission line susceptances, a phenomenon previously described in Bichler & Knörr, (2023).
Figure 2 illustrates nodal IP, CH, and Join prices, averaged over all hours and sorted in ascending order. The dashed horizontal and vertical lines mark the 5th and 95th percentiles for prices and nodes, respectively. Additionally, the figure shows the average national prices and the national average including a uniform EUR/MWh price adder for redispatch costs. This price adder is calculated as the total redispatch costs under min-volume redispatch (ca. EUR 310.74 million) divided by the total demand (ca. 83.3 TWh), and equals 3.73 EUR/MWh. Notably, with redispatch taken into account, higher average prices are expected at some of the nodes. For the majority of nodes, however, the change in average prices is minor or even negative under nodal pricing.
@@
We observe the highest prices in Northern Central Germany in a region with a high share of expensive biomass energy and the lowest prices in the Alps region with inexpensive hydro energy. The associated nodes are relatively isolated with frequent congestion. This behavior might follow from imprecise estimations of generation costs and line limits described in Section 3, yet price outliers can also be observed in the ENTSO-E report (ENTSO-E,, 2022).
@@
A primary goal in the design of new bidding zones is the reduction of price variance within a zone (ACER, 2022b, ), which signals congestion. Tables 5 and 6 indicate that the standard deviation of zonal prices does not decrease with the number of zones. However, varying prices can have two main reasons: congestion and temporal changes in supply and demand. Figure 5 considers hourly standard deviations of prices across all nodes, meaning it illustrates the price variance based on grid congestions in each hour, which we refer to as congestion-based standard deviation. In contrast, Figure 6 includes histograms of the standard deviation of nodal prices across all hours, meaning it illustrates the price variance based on temporal effects of supply and demand at each node, which we refer to as time-based standard deviation. Together, Figures 5 and 6 represent a composition of the standard deviations reported for nodal prices in Table 5. Note that prices at a single node can vary substantially over time due to changes in supply and demand. This time-based standard deviation exceeds the congestion-based standard deviation. For example, IP prices exhibit a mean congestion-based standard deviation of EUR/MWh 12.68, compared to a mean time-based standard deviation of EUR/MWh 17.39. These numbers provide evidence that temporal fluctuations of supply and demand have greater impact on price variance compared to congestion.
@@
Tables 7 and 8 provide a more detailed variance decomposition for IP prices under all zonal configurations. In particular, Table 7 indicates the congestion-based standard deviation of nodal IP prices, averaged over all nodes assigned to a zone. Equivalently, Table 8 summarizes the time-based standard deviations grouped by zone. Specifically, the time-based standard deviation always exceeds the congestion-based standard deviation, implying that temporal effects drive price variance more than congestion effects. At the same time, the time-based standard deviation varies only little across allocation rules, suggesting that nation-wide fluctuations in supply and demand, on average, affect all price zones equally. In contrast, the congestion-based standard deviation has a slightly decreasing trend with more zones, indicating that zonal splits can reduce price variance at least for some of the zones (e.g., Zone 3). This effect, however, is not consistent across all zones, implying that the cross-zonal lines fail to encompass all congested lines. Consequently, the computed dispatch does not account for the remaining intra-zonal congestion, necessitating costly redispatch under all zonal models, as shown in Table 2. In a complementary paper, we demonstrate that the proposed zone configurations lack stability and fail to effectively segregate nodes along congested lines As a result, the congestion-based standard deviation exhibits minimal reduction as the number of zones increases, and redispatch costs remain high.
@@
4.3 MWPs
The average daily MWPs are summarized in Table 9.
@@ 4.3 MWPs
We observe that only the Euphemia algorithm generates zero MWPs, albeit at the expense of welfare losses. An interesting observation is that the welfare losses of Euphemia – as discussed in Section 4.1 – exceed, on average, the MWPs required under a national welfare-maximizing allocation, irrespective of the pricing rule. This implies that the additional generated welfare could potentially cover the MWPs under a national allocation, resulting in a net welfare gain. Under nodal pricing, MWPs tend to increase compared to zonal pricing, particularly for IP pricing. This can be attributed to the fact that under a national allocation rule, there exists a single hourly IP price set at the variable cost of the marginal generator. With positive fixed costs and constant variable costs, such a generator will inevitably incur a loss and necessitate MWPs under IP pricing. In contrast, nodal pricing involves several hourly IP prices and price-setting generators across the network, leading to higher overall MWPs. Conversely, the Join pricing rule generates lower MWPs than IP and CH pricing, constituting a negligible share of the total costs outlined in Table 2. MWPs follow similar patterns across different zonal configurations, but increase with nodal pricing.
4.4 Analysis of Representative Days
The generation costs for each of these days are presented in Table 10. On 2009/11/23, characterized by warm and windy conditions with low demand, generation costs were minimal, and no congestion occurred in the zonal models, resulting in identical allocations. In contrast, February 18, as a cold and cloudy day, saw substantially higher generation costs, but cross-zonal flow constraints had little impact on the allocation. Across all days, the shift from zonal to nodal pricing led to a similar increase in generation costs.
in kEUR National 2 Zones (k) 2 Zones (s) 3 Zones 4 Zones Nodal
2009/11/23 26,673.74 26,673.74 27,088.39 26,673.74 26,818.18 29,596.10
2009/03/10 31,401.65 31,494.31 31,494.38 31,661.73 31,647.02 35,200.69
2009/02/18 46,287.23 46,471.33 46,287,57 46,287.34 46,432,68 49,303.25
Table 10: Generation Costs: Nov 23, Mar 10, Feb 18
in kEUR National 2 Zones (k) 2 Zones (s) 3 Zones 4 Zones Nodal
2009/11/23 5,060.40 5,060.40 4,815.01 5,057.49 4,809.61 0
2009/03/10 5,195.52 5,185.63 5,172.92 5,336.58 5,149.11 0
2009/02/18 4,683.00 4,887.07 4,688.39 4,680.31 4,807.46 0
Table 11: Min-Cost Redispatch Costs: Nov 23, Mar 10, Feb 18
in kEUR National 2 Zones (k) 2 Zones (s) 3 Zones 4 Zones Nodal
2009/11/23 31,734.14 31,734.14 31,903.40 31,731.23 31,627.79 29,596.10
2009/03/10 36,597.17 36,679.94 36,667.30 36,998.31 36,796.13 35,200.69
2009/02/18 50,970.23 51,358.40 50,975.96 50,967.65 51,240.14 49,303.25
Table 12: Total Costs: Nov 25, Mar 10, Feb 18
(a) 2009/11/23 - Avg IP Prices
(b) 2009/11/23 - Avg CH Prices
(c) 2009/11/23 - Avg Join Prices
(d) 2009/03/10 - Avg IP Prices
(e) 2009/03/10 - Avg CH Prices
(f) 2009/03/10 - Avg Join Prices
(g) 2009/02/18 - Avg IP Prices
(h) 2009/02/18 - Avg CH Prices
(i) 2009/02/18 - Avg Join Prices
Figure 7: Average Nodal Prices
@@ 4.5 Discussion
Our numerical experiments with the BZR data set provide insights into different market clearing models, pricing rules, and their short-term economic implications on day-ahead markets. Nodal pricing tends to result in slightly higher average prices compared to zonal pricing, irrespective of the chosen zonal configuration. However, transmission constraints can be violated and necessitate costly redispatch measures. As a result, a zonal pricing selects an inefficient dispatch, and even if redispatch is conducted at minimal costs, the total system costs will be significantly higher than under nodal pricing. Given the steep increase of redispatch costs in recent years, this effect is likely to magnify in the future.
@@
Interestingly, the choice of zonal configuration did not substantially impact prices. All zonal configurations suggested by ACER in the BZR simplify the network to an extent that opting for two, three, or four zones yields similar results regarding the average prices and the price standard deviation. Analyses of individual days show that congestion arises, but not necessarily at the cross-zonal links but within zones. While nodal prices may vary between nodes, these differences incentivize short-term demand response on nodes where this is most helpful for overall system stability. If such price differences persist over more extended time periods, they also set investment incentives for generators and grid expansion. Note that our analysis does not take such incentives into account and assumes static demand. Taking incentives for demand reduction and targeted investments into account would lead to more similar prices rather in the long run.
@@
Given the absence of Walrasian equilibria, any pricing rule – regardless of zonal or nodal pricing – must balance economic properties such as welfare gains and the participants’ lost opportunity costs. The currently used Euphemia algorithm leads to welfare losses and is computationally expensive but avoids paradoxically accepted bids. Our experiments suggest that non-uniform pricing rules achieve low make-whole payments for paradoxically accepted bids while maintaining the optimal allocation and scaling in polynomial time. The findings provide support for non-uniform pricing as it is currently being discussed in the context of the Capacity Allocation and Congestion Management (CACM) Regulations (All NEMO Committee,, 2023). CH pricing minimizes GLOCs, but our results confirm previous findings that high MWPs are necessary. IP pricing, the prevailing rule in U.S. markets, ensures effective congestion signals but requires very high MWPs. The Join pricing rule strikes a balance between LLOCs and MWPs, ensuring low side-payments and good congestion signals.

### v6

@@ 3.1.1 Welfare losses with Euphemia - Impact of Non-Uniform Pricing
Euphemia maximizes welfare subject to linear and uniform prices, which is different from the allocation problem (1). In our analysis, the average daily welfare loss associated with Euphemia compared to the national configuration (without any network constraints) is 0.34% – or EUR 109,801.24 in absolute terms. This is comparable with previous estimate of 0.05% as a relative loss (Meeus et al., , 2009). [^17] Note that the 0.05% relative welfare loss does not correspond to the official Euphemia implementation, but to a simulation that considers block order restrictions (Meeus et al., , 2009). Disregarding market coupling, the cuts introduced by the Euphemia algorithm to get to linear prices do not deteriorate the outcome by a lot. After redispatch, Euphemia suffers an average daily loss of 7.9% (or kEUR 2,859.80) in total costs compared to the nodal allocation rule. We argue that the national, zonal, or Euphemia models differ only marginally in terms of allocation, and only a transition to a nodal allocation rule would substantially affect the average generation and redispatch costs.
@@ 3.2 Price Levels
Table 4 provides average prices over all hours of the weeks under consideration, while Table 5 price standard deviations for all computed prices under the respective allocation and pricing rule. The median prices are presented in D.
@@
Average CH (Convex Hull) prices are slightly lower compared to the other pricing rules under consideration, with lower standard deviation and irrespective of the allocation rule. In contrast, median prices are at comparable levels across all pricing rules. Since CH prices are unaffected by pre-determined generator commitments, they tend to be lower and more stable. In particular, prices can be set by a cheaper – yet uncommitted – generator (Schiro et al., , 2016). In contrast, average Euphemia prices tend to be slightly higher than zonal prices, while median prices are in the same ranges. This finding implies that the price-setting generator is usually identical for Euphemia and zonal prices, yet sometimes the Euphemia algorithm introduces a cut that disallows a generator that would otherwise be dispatched. As a result, a higher-price generator may set the price.
@@
Under nodal pricing (ca. 40,000 prices per day for 1670 nodes – compared to, e.g., 24 national prices per day) more price outliers were observed than under zonal pricing. For instance, 0.13% of IP (Integer Programming) prices had to be capped 100 EUR/MWh under the zonal configurations, compared to 0.25% under nodal pricing. Without capping, the maximum observed IP and Join price would be as high as 62,000 EUR/MWh for an hour of extreme scarcity. Note that SDAC also imposes a price cap of 5,000 EUR/MWh for such periods. CH prices, in contrast, are less prone to outliers, peaking at 423 EUR/MWh. Thus tight flow constraints can increase price levels in the short run. The zonal allocation rules oversimplify the transmission network to an extent that nodal prices can hardly be lower, even without congestion. In contrast, when congestion is present in parts of the grid, nodal prices will be higher, resulting in higher average and median prices. Upon further analysis, it was found that nodal prices have high sensitivity to transmission line susceptances, a phenomenon previously described in Bichler & Knörr, (2023).
Figure 2 illustrates nodal IP, CH, and Join prices, averaged over all hours and sorted in ascending order. The dashed horizontal and vertical lines mark the 5th and 95th percentiles for prices and nodes, respectively. Additionally, the figure shows the average national prices and the national average including a uniform EUR/MWh price adder for redispatch costs. This price adder is calculated as the total redispatch costs (ca. EUR 478 million) divided by the total demand (ca. 83.3 TWh), and equals 5.73 EUR/MWh. Notably, with redispatch taken into account, higher average prices are expected at some of the nodes. For the majority of nodes, however, the change in average prices is minor or even negative under nodal pricing.
@@
We observe the highest prices in Northern Central Germany in a region with a high share of expensive biomass energy and the lowest prices in the Alps region with inexpensive hydro energy. The associated nodes are relatively isolated with frequent congestion. This behavior might follow from imprecise estimations of generation costs and line limits described in Section 2.2, yet price outliers can also be observed in the ENTSO-E report (ENTSO-E, , 2022).
@@
A primary goal in the design of new bidding zones is the reduction of price variance within a zone (ACER, 2022c, ), which signals congestion. Tables 5 and 6 indicate that the standard deviation of zonal prices does not decrease with the number of zones. However, varying prices can have two main reasons: congestion and temporal changes in supply and demand. Figure 5 considers hourly standard deviations of prices across all nodes, meaning it illustrates the price variance based on grid congestions in each hour, which we refer to as congestion-based standard deviation. In contrast, Figure 6 includes histograms of the standard deviation of nodal prices across all hours, meaning it illustrates the price variance based on temporal effects of supply and demand at each node, which we refer to as time-based standard deviation. Together, Figures 5 and 6 represent a composition of the standard deviations reported for nodal prices in Table 5. Note that prices at a single node can vary substantially over time due to changes in supply and demand. This time-based standard deviation exceeds the congestion-based standard deviation. For example, IP prices exhibit a mean congestion-based standard deviation of EUR/MWh 12.68, compared to a mean time-based standard deviation of EUR/MWh 17.39. These numbers provide evidence that temporal fluctuations of supply and demand have greater impact on price variance compared to congestion.
@@
Tables 7 and 8 provide a more detailed variance decomposition for IP prices under all zonal configurations. In particular, Table 7 indicates the congestion-based standard deviation of nodal IP prices, averaged over all nodes assigned to a zone. Equivalently, Table 8 summarizes the time-based standard deviations grouped by zone. Specifically, the time-based standard deviation always exceeds the congestion-based standard deviation, implying that temporal effects drive price variance more than congestion effects. At the same time, the time-based standard deviation varies only little across allocation rules, suggesting that nation-wide fluctuations in supply and demand, on average, affect all price zones equally. In contrast, the congestion-based standard deviation has a slightly decreasing trend with more zones, indicating that zonal splits can reduce price variance at least for some of the zones (e.g., Zone 3). This effect, however, is not consistent across all zones, implying that the cross-zonal lines fail to encompass all congested lines. Consequently, the computed dispatch does not account for the remaining intra-zonal congestion, necessitating costly redispatch under all zonal models, as shown in Table 2. Dobos et al., (2025) demonstrate that the proposed zone configurations lack stability and fail to effectively segregate nodes along congested lines As a result, the congestion-based standard deviation exhibits minimal reduction as the number of zones increases, and redispatch costs remain high.
Before moving to the next section, let us emphasize that day-ahead prices alone cannot fully determine the desirability of different market designs or bidding zone configurations. These prices do not reflect: (1) the magnitude of redispatch costs, (2) network tariffs (Netzentgelte), which TSOs use to cover redispatch costs and maintain network infrastructure, (3) congestion revenues received by TSOs, and (4) payments required to address non-convexities in generation profiles (see Section 2.1.3).
@@
3.3 GLOCs, LLOCs, MWPs
The average daily GLOCs, LLOCs, and MWPs are summarized in Tables 911, respectively. Recall from Section 2.1.3 that GLOCs represent the opportunity costs incurred if agents are allowed to deviate from the current allocation to achieve their individual maximum profit, given the market-clearing prices. LLOCs are a subclass of GLOCs quantifying the opportunity costs obtained when agents are not allowed to change their commitment status. MWPs are the minimum compensation provided to market agents, ensuring they do not incurr losses (i.e., negative utility) given the market-clearing prices. GLOCs and LLOCs are not paid in practice. MWPs, however, are paid in US electricity markets and are generally not covered directly by revenues obtained from the buyers through energy purchases in the day-ahead market.
in EUR National 2 Zones (k) 2 Zones (s) 3 Zones 4 Zones Nodal
IP 2,975,070.52 2,666,810.11 3,332,450.37 2,563,874.88 3,130,778.9 5,469,460.3
CH 124,235.5 162,896.45 520,529.19 141,048.26 717,983.02 382,104.15
Join 981,496.35 1,880,938.84 1,001,437.75 1,801,132.46 2,616,044.17 4,792,706.62
Euphemia 1,002,357.56
Table 9: Average Daily GLOCs
in EUR National 2 Zones (k) 2 Zones (s) 3 Zones 4 Zones Nodal
IP 0.00 0.00 0.00 0.00 0.00 0.00
CH 11,388.11 31,280.46 109,165.0 22,230.84 174,685.4 184,626.79
Join 883.43 3,539.37 1,432.85 1,473.60 15,489.43 36,927.06
Euphemia 0.00
Table 10: Average Daily LLOCs
@@ 3.3 GLOCs, LLOCs, MWPs
By definition, CH prices minimize GLOCs, and on average, they are substantially lower compared to both IP and Join prices, irrespective of the chosen allocation rule. Similar orders of magnitude for GLOCs are observed across all zonal allocation rules, including Euphemia. In contrast, nodal pricing results in the highest GLOCs, aligning with the higher average prices presented in Table 4. Higher prices imply greater forgone profits for generators and increase GLOCs. Generators in nodal markets face stronger incentives to deviate from the optimal outcome, necessitating penalties – a common practice in many U.S. markets (Bichler et al., , 2022).
As a natural subset of GLOCs, LLOCs follow a similar pattern, with the highest values occurring under nodal pricing, once again correlated with higher average prices. IP pricing ensures zero LLOCs, as discussed in Section 2, ensuring reliable congestion signals, meaning that price differences between adjacent nodes arise only if the line connecting them is congested. In this case, the price difference precisely reflects the marginal value of transmission capacity. In contrast, CH pricing consistently results in the highest LLOCs, supporting observations that congestion signals may be flawed (Schiro et al., , 2016). This is especially relevant under a nodal allocation rule considering all transmission lines. The Join pricing rule yields substantially fewer LLOCs than CH pricing, indicating a superior quality of congestion signals.
@@
In terms of average MWPs to compensate generator losses, only the Euphemia algorithm generates zero MWPs, albeit at the expense of welfare losses. An interesting observation is that the welfare losses of Euphemia – as discussed in Section 3.1 – exceed, on average, the MWPs required under a national welfare-maximizing allocation (i.e., a zonal model as described in Section 2.1.2 with a single zone), irrespective of the pricing rule. This implies that the additional generated welfare could potentially cover the MWPs under a national allocation, resulting in a net welfare gain. Under nodal pricing, MWPs tend to increase compared to zonal pricing, particularly for IP pricing. This can be attributed to the fact that under a national allocation rule, there exists a single hourly IP price set at the variable cost of the marginal generator. With positive fixed costs and constant variable costs, such a generator will inevitably incur a loss and necessitate MWPs under IP pricing. In contrast, nodal pricing involves several hourly IP prices and price-setting generators across the network, leading to higher overall MWPs. Conversely, the Join pricing rule generates lower MWPs than IP and CH pricing, constituting a negligible share of the total costs outlined in Table 2.
In summary, GLOCs, LLOCs, and MWPs follow similar patterns across different zonal configurations, but increase with nodal pricing. Concerning pricing rules, IP, CH, and Euphemia each minimize a specific class of lost opportunity costs. However, the Join pricing rule stands out by striking a remarkable balance between low MWPs and LLOCs. At the expense of slightly increased GLOCs, Join prices facilitate marginal side-payments and maintain a high quality of congestion signals.
@@ 3.4 Discussion
Our numerical experiments with the BZR data set provide insights into different market clearing models, pricing rules, and their short-term economic implications on day-ahead markets. Nodal pricing tends to result in slightly higher average prices compared to zonal pricing, irrespective of the chosen zonal configuration. However, in zonal pricing, transmission constraints can be violated and necessitate costly redispatch measures. As a result, a zonal pricing selects an inefficient dispatch, and even if redispatch is conducted at minimal costs, the total system costs will be significantly higher than under nodal pricing. Given the steep increase of redispatch costs in recent years, this effect is likely to magnify in the future.
@@
Interestingly, the choice of zonal configuration did not substantially impact prices. All zonal configurations suggested by ACER in the BZR simplify the network to an extent that opting for two, three, or four zones yields similar results regarding the average prices and the price standard deviation. Analyses of individual days show that congestion arises, but not necessarily at the cross-zonal lines but within zones. While nodal prices may vary between nodes, these differences incentivize short-term demand response on nodes where this is most helpful for overall system stability. If such price differences persist over more extended time periods, they also set investment incentives for generators and grid expansion. Note that our analysis does not take such incentives into account and assumes static demand. Nevertheless, long-run effects related to inefficient placement and technology choice for production, consumption and grid expansion are important factors to consider when evaluating a market clearing model.
@@
Given the absence of Walrasian equilibria, any pricing rule – regardless of zonal or nodal pricing – must balance economic properties such as welfare gains and the participants’ lost opportunity costs. The currently used Euphemia algorithm leads to welfare losses and is computationally expensive. It avoids paradoxically accepted bids but may result in paradoxically rejected bids. Our experiments show that the welfare losses of Euphemia exceed the alternative costs of make-whole payments. Moreover, non-uniform pricing rules, such as CH pricing, achieve low make-whole payments for paradoxically accepted bids while maintaining the optimal allocation and scaling in polynomial time. The findings provide support for non-uniform pricing as it is currently being discussed in the context of the Capacity Allocation and Congestion Management (CACM) Regulations (All NEMO Committee, , 2023). CH pricing minimizes GLOCs, but our results confirm previous findings that high MWPs are necessary. IP pricing, the prevailing rule in U.S. markets, ensures effective congestion signals but requires very high MWPs. The Join pricing rule strikes a balance between LLOCs and MWPs, ensuring low side-payments and good congestion signals.
Finally, we note that a a potential split of the German BZ has raised concerns about long-term regional impacts. Literature shows that energy-intensive industries in the south would likely face higher electricity prices, threatening their competitiveness (Growitsch et al., , 2014; Traber & Kemfert, , 2015). RES providers in the north may benefit from reduced curtailment but could experience lower and more volatile prices, weakening investment incentives (Kunz & Zerrahn, 2015b, ; Hirth & Schlecht, , 2019). A zone split may reduce liquidity in the German futures market, complicating hedging and increasing risk premiums. While such distributional aspects are not the focus of our study, they need to be considered when evaluating the effects of a zonal split.